# The Birth of Be Star Disks II. A High-Resolution Spectroscopic Campaign and TESS Observations of an Outburst of the Classical Be star Pavonis

### v1

@@ Abstract
Be stars are non-supergiant, rapidly rotating B stars that have shown emission lines originating in a circumstellar disk. The exact mechanisms that lead to disk formation and dissipation are not fully known although progress has been made with some systems. Here, we present a study of a disk outburst of the southern Be star Pavonis (HD 173948). Our dataset comprises 698 high‑resolution spectra taken contemporaneously with TESS photometry in 2023. During the final days of TESS monitoring, the star began building a disk from a pristine diskless state. We find that the disk built within 5 days in optical H I and He I lines, while the disk is circularized in about 12 days. The disk began to decay in higher energy He I first, then lower energy transitions, with the decay ending last for H. We examine non-radial pulsations both through TESS photometry and the line profile variations in Balmer lines, He I lines, and the weak photospheric Si III 5739 line. Our analysis indicates that two periodicities seen in TESS photometry (at 1.644 and 1.485 cycles d-1) are not seen in the spectral lines before, during, or after the outburst. The strongest spectral signal is a periodicity at 0.163 cycles d-1, which appears as a difference between the weaker signals. We additionally find evidence for fast non-photometric pulsational variations over the course of spectroscopy obtained before, during, and after the outburst. These fast LPVs are strong, and interfere with the two weaker signals, causing their apparent incoherence.
1 Introduction
@@
Be stars are rapidly rotating, non-supergiant B-type stars that have shown Balmer-line emission arising from a circumstellar Keplerian decretion disk (e.g., Collins_1987; 2013A&ARv..21...69R). The disks are variable on short, observable time scales, which are well described by the Viscous Decretion Disk (VDD) framework (2006ApJ...639.1081C; 2008ApJ...684.1374C; 1991MNRAS.250..432L). One active area of observational research for understanding these stars is correlating various observable signatures such as line equivalent widths (EWs) and the continuum flux with the growth and dissipation of these disks. Recently, 2025arXiv250407571L (hereafter, Paper 1) presented work for 30 disk outbursts that were observed both with spectroscopy and photometrically with the Transiting Exoplanet Survey Satellite (TESS; 2015JATIS...1a4003R). We present in this paper a more in-depth analysis of the Be star Pavonis (HR 7074, HD 173948) that was reported on in Paper 1. Our analysis extends to include five sectors of TESS photometry (Sectors 13, 66, 67, 93, and 94) along with the high-resolution spectra discussed in Paper 1, with the generalized data reduction being described previously in Paper 1. Paper 1 focused primarily on the H behavior and the circularization process and documented the full spectroscopic dataset. We focus on complementary analyses so that in addition to H, we study H, He I 5876, 6678, 4921, 4713 and Si III 5739, with these lines chosen as they are available with both spectrographs used, CHIRON on the CTIO 1.5m (2013PASP..125.1336T) and NRES on the LCO 1.0m (2014SPIE.9147E..16E). Our expanded analysis includes (i) time-resolved EW and V/R (ratio of EW of violet side to EW of red side of the line) fitting for the seven aforementioned optical spectral lines, (ii) an in-depth pulsational study connecting multi-epoch TESS photometry with spectroscopic line-profile variability (LPV), and (iii) a high-cadence intra-night LPV characterization using nine dense sequences. The datasets used are presented in Table 1.
@@
Prior to the presentation of these data and our analysis, we consider a few relevant observational and theoretical considerations related to Be stars and their disks. Several physical processes have been suggested to explain how the disks around Be stars are formed. For example, stellar winds confined to the equatorial region were proposed by 1993ApJ...409..429B to explain the observed disks, but this is not likely the dominant mode for formation (1996ApJ...472L.115O; 1995ApJ...440..308C). Likewise, processes involving strong magnetic fields have been proposed (e.g. 2002ApJ...578..951C), but ultimately rejected (2016ASPC..506..207W; 2018MNRAS.478.3049U). To date, it seems that the most viable explanation can be found in non-radial pulsations (NRPs), as suggested by 1986PASP...98...30O. 2016A&A...588A..56B showed that multiple NRP modes could provide enough energy to form a disk, and 2022AJ....163..226L illustrated that pulsations are very often organized in closely-spaced groups of frequencies. In a specific example, 2020A&A...644A...9N found that certain NRP modes allowed for the transfer of energy towards the surface layers of HD 49330, aiding in an outburst. Furthermore, 2021MNRAS.508.2002R demonstrated that a periodic phenomenon of disk building for HD 6226 was related to two closely-spaced NRP frequencies. Other earlier works provide a theoretical framework for NRPs (e.g. 2009ApJ...701..396C), and observational evidence of NRPs was presented for the star Centauri by 1998ASPC..135..343R.
@@
Disk formation and dissipation for Be stars can aid our understanding of astrophysical disks, and thus observations are needed to constrain how the disks evolve. We can measure the changing disk through the strength of emission lines, which generally increase during disk growth, and decrease during dissipation. The long-term viscous evolution of the disk is well described by the Viscous Decretion Disk (VDD) framework (2006ApJ...639.1081C; 2008ApJ...684.1374C; 1991MNRAS.250..432L). Paper 1 discusses the circularization and VDD modeling in detail, so we adopt that framework and focus our new analyses on multi-line EW and V/R evolution.
@@
| Sector | Date range (civil) | HJD | # points | Cadence (min) | Noise Floor |
@@
| 66–67 | 2023 Jun 02 – 2023 Jul 29 | 2460097 - 2460155 | 22,027 | 2 | 0.0054 |
| 93–94ddPhotometric data not previously analyzed in Paper 1. | 2025 Jun 03 – 2025 Jul 25 | 2460829 - 2460882 | 20,417 | 2 | 0.0054 |
This paper focuses on the southern star Pav, a member of the Scorpio-Centaurus association, with spectral type B2IIIe (10.1111/j.1365-2966.2006.10655.x). Early works on Pav (e.g. 1949Obs....69...30G; 1950ApJ...111..663S) focused on describing emission patterns. Later, 1975ApJS...29..137S and 1989ApJ...347.1082C measured rotational velocities and found them to be either km s-1 or km s-1. Most recently, 2016A&A...595A.132Z measured a of 145 km s-1, which we use later in this work. Since these earlier studies, the focus has been on studying the variability of different emission lines. 2003A&A...411..229R noted that Pav has line profile variability caused by NRPs, but lacked enough data to determine a period. 2011A&A...533A..75L performed Fourier analysis on time series spectroscopy, finding four frequencies: = 0.17 0.02, = 0.49 0.05, = 0.82 0.03, and = 1.63 0.04 cycles d-1, which were interpreted as g-mode pulsation (), orbiting circumstellar material (), and likely aliases (i.e. without a physical explanation; and ). Paper 1 found that this was a system that began ejecting material to form a disk at the end of two consecutive TESS sectors, which we further explore here to include additional spectral lines and pulsations in this paper. Paper 1 further discusses the observational history and their findings. Notably, 2011A&A...533A..75L found a frequency of 0.82 d-1 that was interpreted as being circumstellar in origin, which is also seen in the emission oscillation asymmetry cycles (the V/R frequency/period during the early stage of the outburst) derived in Paper 1. But this frequency is also present in TESS photometry even when there is no disk material, as evident by the lack of any emission in contemporaneous spectroscopy. As these results are the main findings related to Pav in Paper 1, an in-depth analysis may be able to provide a better understanding of how the disk forms around this Be star, especially since this disk growth happened at a time when when there was no no hint of any pre-existing circumstellar material.
@@
2 A Well-Observed Disk Outburst in 2023
the line profile variations, especially on the dominant d period, are seen as darker perturbations moving from blue t red.
@@ 2 A Well-Observed Disk Outburst in 2023
It is worth understanding the original observing strategy for the spectrscopic data collection in order to put our results in context. Our observational campaign began with the hope to understand the spectroscopic and photometric properties of Pav’s NRPs as the star showed no sign of a disk in the early times we observed the star with NRES during TESS Sector 66. In case of an outburst, we monitored the star and checked H for emission with NRES with additional intense monitoring for pulsational signatures and fast variations from CHIRON during the TESS observation window. For the majority of this time, no emission features were present. We then observed the sudden appearance of H and H emission on 2023 July 22 (HJD 2460148), roughly a week prior to the end of the TESS observations. We changed our approach for observing the target, continuing the spectroscopic monitoring after the TESS observations and obtaining additional intense time-series to examine potential pulsational properties both during and after an outburst of disk activity in this Be star.
@@
In the top panels of Fig. 1 are spectral line profiles for seven lines that were used in this study available in both the CHIRON and NRES spectra. We highlight spectra from before, during, and after the outburst. The bottom panels of Fig. 1 show the evolution of H and He I 5876 lines in a dynamical representation and indicates that emission began around July 26, 2023 (HJD 2460152). There was a gap in observations around this date, so what we can say definitively is that there were no signs of emission before July 26, but that emission had started afterwards in all lines. This agrees with the EW measurements, which indicate that emission started concurrently in all lines with emission. By 2023 August 10 (HJD 2460166), emission levels had dissipated in all He I lines, but Balmer emission levels had not dissipated to the continuum level until 2023 August 20 (HJD 2460177). The dynamical representation displays the fast buildup of the disk with the material in emission that is seen to move around the profile center, until the line is more symmetric as the disk circularized, and finally begins to dissipate. Furthermore, especially in the He I profile, the difference profiles in Fig. 1 show clear d blue-to-red perturbations, which is most easily seen outside of the disk outburst.
@@
In addition to the visual inspection, we measured the EW for the spectral lines displayed in Fig. 1. In Fig. 2, we compare EW measurements for H, H, He I 4713, He I 4921, He I 5875, He I 6678, and Si III 5739 with the TESS flux where we indicate the timing of the outburst for each set of measurements in gray. The Balmer and He I lines rise sharply in their emission beginning around July 22/July 27 (HJD 2460148/2460153), corresponding to the start of emission and the peak in flux in the TESS data. Then, EW values of the He I lines gradually return to pre-outburst levels by August 10 (HJD 2460166), consistent with the dissipation of the disk, and the end of the disk growth. The higher excitation He I lines at 4713 and 4921 Å returned to their pre-outburst level faster than the He I 5875 and 6678 lines. Outside of the outburst window, the EW curves are stable, confirming that Pav underwent the entire outburst event completely contained within a roughly three-week time frame. However, we note that there may be a slight rise in the level of the line emission following the outburst compared to prior to the outburst.
@@
With the data we have available to us, we can quantify the decay of the disk. The VDD model is described by 2006ApJ...639.1081C; 2008ApJ...684.1374C, and has had success in many cases (e.g 2021MNRAS.508.2002R, and references therein). 2016ASPC..506..157R found a general form for the disk dissipation given by
@@
In this equation, is the time-dependent magnitude, is the quiescent magnitude, is the amplitude of the decay, dictates the timescale of the decay of the light curve, and is a zero-point in time. The VDD model has been used for the decay of the disk, but a similar model for growth has not yet been established for short outbursts like the one we observed for Pav, though 2018MNRAS.476.3555R have developed models for longer-duration outbursts. This exponential decay has been used previously to describe the decay of the disk signatures in spectral lines (e.g. 2021MNRAS.508.2002R), so we use a form that handles our EW values, namely
@@
We can fit the hydrogen and He I lines with an MCMC sampler procedure using Eq. 2 as our model, along with the data near the peak of the outburst and onwards to the end of the 2023 spectroscopic time-series data. We developed the model with the emcee code described by 2013PASP..125..306F, which ran with 256 walkers, and after running a 100‑step burn‑in, was followed by 1,024 iterations. Our posterior spread of models was examined with the corner package in python (corner) indicating that high-quality fits were made, with an example shown at the end of this paper in Fig. 7. The maximum likelihood model is also calculated via the lnprob() function and is the posterior sample with the largest recorded log-likelihood after burn-in. The Si III 5739 line from Fig. 2 shows no detectable emission, so it may not be useful for understanding the disk, but various lines that contain no emission will be useful for understanding pulsational behavior. With the fits are their associated posterior spreads in Fig. 2 on top of the measurements. From the fits, we include in Table 2 the parameters and the errors given in the Monte Carlo sampling output. We see that the hydrogen lines decay over a longer time frame (=8.5 days for H and =6.4 days for H) than the helium lines ( 2.4–5.6 days). With a positive exponent rather than a negative, we fit the rise of the disk with results in Table 2 and Fig. 2, although this is not as physically motivated as the decay parameters. Within Table 2, not all errors are symmetric, and may be underestimated (see e.g. 2021MNRAS.502.5038N). While these errors are slightly increased for parameters that had an error of 0, the posterior spread in the models in Fig. 2 imply that our fits are realistic.
@@
In Paper 1’s recent analysis of many H and photometric outbursts recorded with TESS, it was demonstrated that V/R variations in Be stars can be described with a time–dependent sinusoid of the form
@@
We fit the resulting V/R time series using an emcee sampler (2013PASP..125..306F) with 240 walkers, running a 100‑step burn‑in followed by 1,024 iterations. Fig. 3 displays the V/R data and our maximum‑likelihood fit with the resulting posterior spread. We also include the corner plots online with the plots from the decay sampler fits. The fits in Fig. 3 illustrate how well constrained the posterior distributions are for V/R0, A, P, , and for all of the spectral lines examined, with the parameters included in Table 2. Although the form used in this work differs slightly from Paper 1 in how it allows amplitude (and period) to vary in that we use an exponential decay compared to a linear decay, the inclusion of ensures a physically meaningful decay in line with the disk’s evolution. The period (1.25 0.01 d) of the sinusoidal term agrees with the H measurements of Paper 1 who reported a frequency of 0.81 c d-1. These fits and the associated error estimates are given in Table 2.
3 Periodic Behavior

### v2

@@ Abstract
Be stars are rapidly-rotating B stars that have shown emission lines originating in a circumstellar disk. The mechanisms that lead to disk formation and dissipation are not known although progress has been made with some systems. We present a study of a disk outburst of the Be star Pavonis. Our dataset comprises 698 high‑resolution spectra contemporaneous with TESS photometry in 2023. Near the end of TESS monitoring, the star began disk building from a pristine diskless state. We find the disk built within 5 days in optical H I and He I lines, while the disk circularized in about 12 days. The disk began to decay in higher excitation He I first, then lower excitation transitions, with the decay ending last for H. We examine non-radial pulsations both through TESS photometry and the line profile variations (LPVs) in the spectroscopy. Our analysis indicates that two periodicities seen in TESS photometry (at 1.644 and 1.485 cycles d-1) are not seen in the spectral lines before, during, or after the outburst. The strongest photometric signal is a periodicity at 0.163 cycles d-1, which appears as a difference between the two weaker signals and is visible in the spectra without any apparent changes in amplitude or phase. We additionally find evidence for fast non-photometric pulsational variations over the course of spectroscopy obtained before, during, and after the outburst. These fast LPVs are strong, and interfere with the two weaker signals, hampering our ability to detect them in spectroscopy.
I Introduction
@@
Be stars are rapidly rotating, non-supergiant B-type stars that have shown Balmer-line emission arising from a circumstellar Keplerian decretion disk (e.g., Collins, 1987; Rivinius et al., 2013). The disks are variable on short, observable time scales, which are well described by the Viscous Decretion Disk (VDD) framework (Carciofi and Bjorkman, 2006, 2008; Lee et al., 1991). One active area of observational research for understanding these stars is correlating various observable signatures such as line equivalent widths (EWs) and the continuum flux with the growth and dissipation of these disks. Recently, Labadie-Bartz et al. (2025) (hereafter, Paper I) presented work for 30 disk outbursts in 13 stars that were observed both with spectroscopy and photometrically with the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015). We present in this paper a more in-depth analysis of the Be star Pavonis (HR 7074, HD 173948) that was reported on in Paper I. Our analysis extends to include five sectors of TESS photometry (sectors 13, 66, 67, 93, and 94) along with the high-resolution spectra discussed in Paper I, with the generalized data reduction being described previously in Paper I. Paper I focused primarily on the H behavior and the circularization111The term circularization may be ambiguous in general. In this work, we use this term to refer to the circumstellar material becoming azimuthally homogenized – i.e. forming an axi-symmetric disk. We are not attempting to describe the early-time orbital paths of individual circumstellar particles, which may have elliptical or circular orbits. process and documented the full spectroscopic dataset. In this paper, we focus on complementary analyses so that in addition to H, we study H, He I 5876, 6678, 4921, 4713 and Si III 5739, with these lines chosen as they are available with both spectrographs used, CHIRON on the Cerro Tololo Inter-American Observatory (CTIO) 1.5-m (Tokovinin et al., 2013) and NRES on the Las Cumbres Observatory (LCO) 1.0-m (Eastman et al., 2014). Our expanded analysis includes (i) time-resolved EW and V/R (ratio of EW of violet side to EW of red side of the line) fitting for the seven aforementioned optical spectral lines, (ii) a pulsational study connecting multi-epoch TESS photometry with spectroscopic line-profile variability (LPV), and (iii) a high-cadence intra-night LPV characterization using nine dense sequences. The datasets used are presented in Table 1.
@@
Prior to the presentation of these data and our analysis, we review a few relevant observational and theoretical considerations related to Be stars and their disks. Several physical processes have been suggested to elucidate how the disks around Be stars are formed. For example, stellar winds confined to the equatorial region were proposed by Bjorkman and Cassinelli (1993) to explain the observed disks, but this is not likely the dominant mode for formation (Owocki et al., 1996; Cranmer and Owocki, 1995). Likewise, processes involving strong magnetic fields have been raised (e.g., Cassinelli et al., 2002), but ultimately rejected (Wade et al., 2016; ud-Doula et al., 2018). To date, it seems that the most viable explanation can be found in non-radial pulsations (NRPs), as suggested by Osaki (1986). Baade et al. (2016) showed that the combination of multiple NRP modes could remove energy and angular momentum deficiencies that limit outbursts, and Labadie-Bartz et al. (2022) illustrated that pulsations are very often organized in closely-spaced groups of frequencies. In a specific example, Neiner et al. (2020) found that certain NRP modes allowed for the transfer of energy towards the surface layers of HD 49330, aiding in an outburst. Furthermore, Richardson et al. (2021) demonstrated that a periodic phenomenon of disk building for HD 6226 was related to two closely-spaced NRP frequencies. On the other hand, or perhaps in parallel, it has been suggested that Rossby waves may become mechanically excited and/or enhanced in visibility during mass ejection activity, perhaps in response to the outburst (Saio et al., 2018; Saio, 2024). Other earlier works provide a theoretical framework for NRPs (e.g., Cranmer, 2009), and observational evidence of NRPs was presented for the star Centauri by Rivinius et al. (1998).
@@
Disk formation and dissipation for Be stars can aid our understanding of astrophysical disks, and thus observations are needed to constrain how the disks evolve. We can measure the changing disk through the strength of emission lines, which generally increase during disk growth, and decrease during dissipation. Paper I discusses the circularization and VDD modeling in detail, so we adopt that framework and focus our new analyses on multi-line EW and V/R evolution.
@@
| Sector | Date range (civil) | HJD | # points | Cadence (min) | Noise FloorddPhotometric data not previously analyzed in Paper I. (ppt) |
@@
| 66–67 | 2023 Jun 02 – 2023 Jul 29 | 2460097 - 2460155 | 22,027 | 3.33 | 0.0054 |
| 93–94eefootnotemark: | 2025 Jun 03 – 2025 Jul 25 | 2460829 - 2460882 | 20,417 | 3.33l | 0.0054 |
This work studies the southern star Pav, a member of the Scorpio-Centaurus association, with spectral type B2IIIe (Levenhagen and Leister, 2006). Early works on Pav (e.g., Gascoigne, 1949; Sahade, 1950) described emission patterns. Later, Slettebak et al. (1975) and Chen et al. (1989) measured rotational velocities and found them to be either km s-1 or km s-1. Most recently, Zorec et al. (2016) measured a of 145 km s-1, which we use later in this work. Since these earlier studies, the focus has been on studying the variability of different lines. Rivinius et al. (2003) noted that Pav has line profile variability caused by NRPs, but lacked enough data to determine a period. Levenhagen et al. (2011) performed Fourier analysis on time series spectroscopy, finding four frequencies: = 0.17 0.02, = 0.49 0.05, = 0.82 0.03, and = 1.63 0.04 cycles d-1, which were interpreted as g-mode pulsation (), orbiting circumstellar material (), and likely aliases (i.e. without a physical explanation; and ). Paper I found that this was a system that began ejecting material to form a disk at the end of TESS sectors 66 and 67, which we further explore here to include additional spectral lines and pulsations in this paper. Paper I additionally reported that also appears in the emission oscillation asymmetry cycles (the V/R frequency/period during the early stage of the outburst). But this frequency is also present in TESS photometry even when there is no disk material, as evident by the lack of any emission in contemporaneous spectroscopy. As these results are the main findings related to Pav in Paper I, an in-depth analysis may be able to provide a better understanding of how the disk forms around this Be star, especially since this disk growth happened at a time when when there was no hint of any pre-existing circumstellar material.
@@
II A Well-Observed Disk Outburst in 2023
@@
It is worth understanding the original observing strategy for the spectroscopic data collection in order to put our results in context. Our observational campaign began with the hope to understand the spectroscopic and photometric properties of Pav’s NRPs as the star showed no sign of a disk in the early times we observed the star with NRES during TESS sector 66. In case of an outburst during TESS sector 67, we monitored the star and checked H for emission with NRES, with additional intense monitoring for pulsational signatures and fast variations from CHIRON during the TESS observation window. For the majority of this time, no emission features were present. We then observed the sudden appearance of H and H emission on 2023 July 26 (HJD 2460152), not long before the end of the TESS observations of Pavonis for the year. We changed our approach for observing the target, continuing the spectroscopic monitoring after the TESS observations and obtaining additional intense time-series to examine potential pulsational properties both during and after the outburst of disk activity in this Be star.
@@
In the top panels of Fig. 1 are spectral line profiles for seven lines that were used in this study available in both the CHIRON and NRES spectra. We highlight spectra from before, during, and after the outburst. The bottom panels of Fig. 1 show the evolution of H and He I 5876 lines in a dynamical representation and indicate that emission began around July 26, 2023 (HJD 2460152). There was a gap in observations around this date, so what we can say definitively is that there were no signs of emission before July 26, but that emission had started afterwards in all lines. This agrees with the EW measurements, which indicate that emission started concurrently in all lines with emission. By 2023 August 10 (HJD 2460166), emission levels had dissipated in all He I lines, but Balmer emission levels had not dissipated to the continuum level until 2023 August 20 (HJD 2460177). The dynamical representation displays the fast buildup of the disk with the material in emission that is seen to move around the profile center, until the line is more symmetric as the disk circularized, and finally begins to dissipate. Furthermore, especially in the He I profile, the difference profiles in Fig. 1 show clear d blue-to-red migrations, which is most easily seen outside of the disk outburst.
@@
In addition to the visual inspection, we measured the EW for the spectral lines displayed in Fig. 1. In Fig. 2, we compare EW measurements for H, H, He I 4713, He I 4921, He I 5875, He I 6678, and Si III 5739 with the TESS flux where we indicate the timing of the outburst for each set of measurements in gray. The Balmer and He I lines rise sharply in their emission beginning around July 26 (HJD 2460152), corresponding to the start of emission and the increase in flux in the TESS data. Then, EW values of the He I lines gradually return to pre-outburst levels by August 10 (HJD 2460166), consistent with the dissipation of the disk, and the end of the disk growth. The higher excitation He I lines at 4713 and 4921 Å returned to their pre-outburst level faster than the He I 5875 and 6678 lines. Outside of the outburst window, the EW curves are stable, confirming that Pav underwent the entire outburst event completely contained within a roughly three-week time frame. However, we note that there may be a slight rise in the level of the line emission following the outburst compared to prior to the outburst.
@@
With the data we have available to us, we can quantify the decay of the disk. The VDD model is described by Carciofi and Bjorkman (2006, 2008), and has had success in many cases (e.g., Richardson et al., 2021, and references therein). Rímulo et al. (2016) found a general form for the disk dissipation given by
@@
In this equation, is the time-dependent magnitude, is the quiescent magnitude, is the amplitude of the decay, dictates the timescale of the decay of the light curve, and is a zero-point in time. These parameters are given in equation 5 in Rímulo et al. (2016), but characterized with filters for Johnson photometry rather than the generalized bandpass we used here. The VDD model has been used for the decay of the disk, but a similar model for growth has not yet been established for short outbursts like the one we observed for Pav, though Rímulo et al. (2018) have developed models for longer-duration outbursts. This exponential decay has been used previously to describe the decay of the disk signatures in spectral lines (e.g., Richardson et al., 2021), so we use a form that handles our EW values, namely
@@
We can fit the hydrogen and He I lines with an MCMC sampler procedure using Eq. 2 as our model, along with the data near the peak of the outburst and onwards to the end of the 2023 spectroscopic time-series data. We developed the model with the emcee code described by Foreman-Mackey et al. (2013), which ran with 256 walkers, and after running a 100‑step burn‑in, was followed by 1,024 iterations. Our posterior spread of models was examined with the corner package in python (Foreman-Mackey, 2016) indicating that high-quality fits were made, with an example shown at the end of this paper in Fig. 7. The maximum likelihood model is also calculated via the lnprob() function and is the posterior sample with the largest recorded log-likelihood after burn-in. The Si III 5739 line from Fig. 2 shows no detectable emission, so it may not be useful for understanding the disk, but various lines that contain no emission will be useful for understanding pulsational behavior. posterior spreads are overplotted with their fits in Fig. 2. From the fits, we include in Table 2 the parameters and the errors given in the Monte Carlo sampling output. We see that the hydrogen lines decay over a longer time frame (=8.5 days for H and =6.4 days for H) than the helium lines ( 2.4–5.6 days). With a positive exponent rather than a negative, we fit the rise of the disk with results in Table 2 and Fig. 2, although this is not as physically motivated as the decay parameters. Within Table 2, not all errors are symmetric, and may be underestimated (see, e.g., Nazé et al., 2021). While these errors are slightly increased for parameters that had an error of 0, the posterior spread in the models in Fig. 2 implies that our fits are realistic.
@@
In Paper I’s recent analysis of many H and photometric outbursts recorded with TESS, it was demonstrated that V/R variations in Be stars can be described with a time–dependent sinusoid of the form
@@
We fit the resulting V/R time series using an emcee sampler (Foreman-Mackey et al., 2013) with 240 walkers, running a 100‑step burn‑in followed by 1,024 iterations. Fig. 3 displays the V/R data and our maximum‑likelihood fit with the resulting posterior spread. We also include the corner plots online with the plots from the decay sampler fits. The fits in Fig. 3 illustrate how well constrained the posterior distributions are for V/R0, A, P, , and for all of the spectral lines examined, with the parameters included in Table 2. Although the form used in this work differs slightly from Paper I in how it allows amplitude (and period) to vary in that we use an exponential decay compared to a linear decay, the inclusion of ensures a physically meaningful decay in line with the disk’s evolution. The period (1.25 0.01 d) of the sinusoidal term agrees with the H measurements of Paper I which reported a frequency of 0.81 c d-1. These fits and the associated error estimates are given in Table 2.
III Periodic Behavior